# Entanglement cones and horizons in analogue cosmological production of Dirac fermions

### v1

@@ Abstract
Given that the Hubble rate $\mathsf{H}$ allows us to smoothly interpolate between sudden and adiabatic limits, it is very natural to extend this EE quasi-particle picture to our cosmological scenario. Moreover, resorting to the concept of entanglement contour (EC) [28], which assigns to each site $i$ of the partition a specific contribution $S_{i}$ to the overall EE, such that $\sum_{i}S_{i}=S_{A}$, one can get further information on the local distribution of entanglement within the block, allowing us to bring this quasi-particle picture to a new domain. In particular, we show below that the EC allows identifying a cosmological particle horizon due to the accelerating spacetime. Although a general recipe for the calculation of EC in generic setups is lacking, the fGS ansatz provides a well-defined approximation to the EC via the aforementioned correlation matrix $\Gamma(\eta)$, as the spectrum of the block density matrix $\rho_{A}$ is determined by the restricted correlation matrix $(\Gamma_{A}(\eta))_{ij}=\Gamma_{ij}(\eta)|_{i,j\in A}$ [51, 52]. Diagonalizing this matrix by a unitary transformation $U^{\dagger}\Gamma_{A}U={\rm diag}(\{\nu_{\textrm{k}}\}_{\textrm{k}\in{\rm BZ}})$, the EC $S_{i}(\eta)=\sum_{\textrm{k}\in{\rm BZ}}p_{i}(\textrm{k},\eta)s_{\textrm{k}}(\eta)$ weighs each contribution $s_{\textrm{k}}(\eta)=-\nu_{\textrm{k}}(\eta)\log\nu_{\textrm{k}}(\eta)-(1-\nu_{\textrm{k}}(\eta))\log(1-\nu_{\textrm{k}}(\eta))$ at the $i$-th site by $p_{i}(k,\eta)=|U_{\textrm{k}i}(\eta)|^{2}$. Since this matrix is determined self-consistently (3), we can explore how entanglement evolves in the spacetime coordinates under interaction effects beyond perturbation theory, e.g. symmetry breaking. Let us start by studying how the particles created during the expansion propagate in the asymptotically-flat out region $\mathsf{a}_{\rm f}$, where the conformal time becomes linearly related to the cosmological time $\eta=\eta_{\rm f}+(t-t_{\rm f})/\mathsf{a}_{\rm f}$ and can attain arbitrarily-large values. In Fig. 1(a), we depict the spatiotemporal EC profile for a partition of $\ell_{A}=128$ sites in a $N_{\rm S}=512$ lattice in real time. In the free case $g_{0}^{2}=0$, one observes a clear light-cone-like propagation of entanglement. In the scaling limit, $\ell_{A}$, $t\to\infty$ with $t/\ell_{A}$ finite, this propagation of EC can be obtained by extending the EE quasi-particle picture to the EC, as discussed in the Supplemental Material [53]. The cones emanate from the block boundaries, separating causally connected and disconnected spacetime regions, and show a slope that is set by twice the group velocity $v_{\rm g}=\max_{\textrm{k}\in{\rm BZ}}\left|\partial_{\textrm{k}}\epsilon_{\textrm{k}}\right|$, where $\epsilon_{\textrm{k}}$ is the energy of a particle or antiparticle obtained by diagonalising $h_{\rm k}$ (1). Interestingly, this picture predicts that the entanglement equilibrates and there is an equipartition across the block $S_{i}(\eta)=S_{0}\,\,\forall i\in A$, such that the EE fulfills a volume law that agrees with a the eigenstate thermalization hypothesis [54, 55, 56] and a generalised Gibbs ensemble (GGE) that approximates $\rho_{A}$ [57].
@@
When we switch on interactions to $g_{0}^{2}=1$, the scalar condensate attains a non-zero value in the initial state and displays a non-trivial time evolution $\Sigma(\eta)$. This results in a new dispersion relation from Eq. (3) that is corrected by a time-dependent self-energy and can, consequently, lead to a dynamical group velocity $v_{\rm g}(\eta)$. In this case, the previous conserved quantities associated to the GGE become time-dependent, and may compromise the validity of the quasi-particle prediction. However, for certain regimes in which the condensate relaxes to a certain value, one expects to be able to derive a hydrodynamic description for the EC by renormalising the spread velocity and the particle-antiparticle mode entanglement. We confirm this expectation in the right panel of Fig. 1(a), where we show that the EC light cones follow $x(\eta)\approx\pm 2{v}^{\rm int}_{g}\eta$, where ${v^{\rm int}_{g}}$ is a renormalised velocity. As shown in the figure, the cones get tilted with respect to $g_{0}^{2}=0$, showing that this propagation speed is enhanced by interactions. Although this is counterintuitive at first, as the continuum QFT predicts a dynamical mass generation that would slow the dynamics of excitations, the problem on the lattice is richer as this dynamically-generated mass can bring one closer or further apart from a critical line around which the continuum limit should be valid. In this case, increasing interactions brings us closer, reducing the gap and also the overall bare mass, such that the particle-anti-particle pairs propagate faster. Let us emphasise that this renormalised quasi-particle picture can actually be completely broken for stronger interactions in which the pseudo-scalar condensate is also non-vanishing $\Pi(\eta)$. We find in that case that the group velocity develops persistent oscillations that make a quasi-particle prediction very inaccurate EC (see the Supplemental Material [53]).
@@
Correlation horizons.– Let us now turn to a key phenomenon that becomes apparent via the EC dynamics during the accelerated de Sitter expansion $\mathsf{a}(\eta)$, and after it as above. For this cosmological expansion, a particle horizon determines when the distance between the emission of a particle and its subsequent detection starts to expand faster than the speed of light. Accordingly, spacetime breaks in causally disconnected regions that can have no influence on one another and, thus, no entanglement. In conformal time, the appearance of this horizon in the EC is a bit subtle. In Fig. 1(b) we show the evolution of EC during the expansion, which shows a ballistic spread as before, but this time is upper bounded by $\eta=0$, as cosmological time squeezes in the distant future as a result of the exponential scale factor  [^1] To ensure significant propagation since the beginning of the expansion, we impose an initial state with matter content. This is achieved via a quantum quench in the mass before the start of the expansion. Other alternatives to design the initial state would lead to similar interplay with the horizon..
@@
In Fig. 1(c), we show the EC profile in cosmological time $t$, which clearly shows how the propagation cones get curved as a consequence of the expansion of the universe, leading to a pair of asymptotes that that separate regions that cannot be causally connected. We see two separated spatial regions where entanglement attains non-vanishing values as time elapses but, due to the accelerated expansion, these are separated by a finite region with a width $\Delta x=\ell_{A}-\frac{4v_{g}}{{\sf H}\mathsf{a}_{0}}$ in which the EC is always vanishing. We emphasize that the spatial resolution brought by the EC as opposed to the EE is crucial to identify these correlation horizons. It is also interesting to note that the EC distribution within the causally-connected parts does not fulfill an equi-partition, so the evolution in this case is very different from an equilibration to a GGE. Had we let the system evolve in the Minkowski out region $\mathsf{a}_{\rm f}$, the particle-anti-particle pairs would have had sufficient time to propagate in this static background, such that the causally-disconnected regions would again be connected by subsequent correlation cones as before, leading to volume-law entanglement and GGE.
@@
Symmetry breaking in the entanglement contour.– We have previously mentioned in passing that expansions in which the pseudo-scalar condensate is non-vanishing $\Pi(\eta)$ can have a bigger impact on the spread of entanglement. We now explore this further, considering that parity gets spontaneously broken, which will impact the properties of the EC under point-like transformations [28]. Specifically, the EC satisfies that $S_{i}=S_{j}$ if the system is invariant under a transformation that exchanges two sites, $\hat{U}_{ij}\hat{\psi}_{i}\hat{U}_{ij}^{-1}=\mathord{\rm e}^{\mathord{\rm i}\xi}\hat{\psi}_{j}$, where $\xi\in\mathbb{R}$ and $i\neq j$. To understand this property better, we temporarily focus on the discrete symmetries of the static model $\mathsf{a}(\eta)=1$, which connects to the tenfold classification of topological insulators [59, 60, 61]. For $\Pi=0$, the system is invariant under time-reversal $\hat{T}$ and particle-hole $\hat{C}$ transformations, as well as their combination $\hat{S}=\hat{T}\hat{C}$. At the single-particle level, these transformations are realised by $\mathsf{T}=T\mathcal{K}$ with $T=\gamma^{0}$, $\mathsf{C}=C\mathcal{K}$ with $C=\gamma^{0}\gamma^{1}$, and $\mathsf{S}=\gamma^{0}\gamma^{5}$, where $\mathcal{K}$ denotes complex conjugation such that $T^{\dagger}h_{-\textrm{k}}^{*}T=h_{\textrm{k}}$, $C^{\dagger}h_{-\textrm{k}}^{*}C=-h_{\textrm{k}}$ and $\mathsf{S}^{\dagger}h_{\textrm{k}}\mathsf{S}=-h_{\textrm{k}}$. These operators satisfy $\mathsf{T^{2}=C^{2}=S^{2}}=1$. In addition, parity symmetry $\hat{P}$ is realised as $P=\gamma^{0}$ such that $P^{\dagger}h_{-\textrm{k}}P=h_{\textrm{k}}$. Indeed, together with translational invariance, parity symmetry describes the invariance under a mirror symmetry about any point and, in particular, about the center of the $A$ block, which leads to a site exchange $\hat{U}_{ij}$ with $j=\ell_{A}+1-i$. According to our previous comment, if parity is conserved, one expects the EC to display the symmetry $S_{i}=S_{\ell_{A}+1-i}$, which agrees with the EC cones of Fig. 1.
@@
This situation changes when $\Pi\neq 0$, as one finds that $\hat{T}$ is preserved, but $\hat{P}$, $\hat{C}$ and $\hat{S}$ get broken. Moreover, switching back to $\mathsf{a}(\eta)\neq 1$ will also break $\hat{T}$, endowing the condensate with its own dynamics $\Pi(\eta)$, which shall affect the EC spreading as we now discuss. In the left panel Fig. 2(a), we depict the spread of the EC in complete analogy to Fig. 1, but for an expansion with a non-zero $\Pi(\eta)$. We note that in all of the case above, the ECs displayed actually correspond to one of the spinor components, namely $S_{i}^{\rm u}(\eta)$, which does not compromise the generality as $S_{i}^{\rm d}(\eta)=S_{i}^{\rm u}(\eta)=:S_{i}(\eta)$. This no longer holds when $\Pi(\eta)\neq 0$, and one can observe in the right panel of Fig. 2(a) that the EC for the down spinor $S_{i}^{\rm d}(\eta)$ is indeed different, with a larger spread of entanglement from the left boundary with respect to the right one , opposite to what happens for $S_{i}^{\rm d}(\eta)$. These figures suggest that the EC profiles are still symmetric by simultaneously exchanging left-right boundaries and up-down spinors, yielding $S^{\textrm{u}}_{i}(\eta)=S^{\textrm{d}}_{\ell_{A}+1-i}(\eta)$ and viceversa $S^{\textrm{d}}_{i}(\eta)=S^{\textrm{u}}_{\ell_{A}+1-i}(\eta)$. In fact, one can check that this is indeed a remnant symmetry that combines both parity $\hat{P}$ and particle-hole $\hat{C}$ symmetries $(\gamma^{1})^{\dagger}h_{\textrm{k}}^{*}\gamma^{1}=-h_{\textrm{k}}$, which still holds even if the individual symmetries are broken. The ${C}{P}$ symmetry that survives a non-zero $\Pi(\eta)$ suggests that one can define a zigzag ordering of the spatial sites and spinor components and depict the site- and spinor-resolved EC as shown in Fig. 2(b). This symmetry-aware representation allows to recover a clear light-cone-like structure characteristic of causality and connecting with the aforementioned quasi-particle picture, even in presence of a pseudo-scalar condensate and for any Hubble rate $\mathsf{H}$.
@@
Figure 2: Time evolution of EC for different phases for a partition $A$ of $N_{A}=32$ sites, from a system of $N_{S}=128$ sites. (a) Numerical time evolution of EC for each leg of the ladder, for an expansion with $\Pi\neq 0$, with $ma=-1$, $g^{2}=0$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $\mathsf{H}=100$ (quench limit). (b) Numerical time evolution of EC, with all the sites shown in the horizontal axis, alternating up and down spinors, for an expansion with $\Pi\neq 0$, with $ma=-1$, $g^{2}=0$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $\mathsf{H}=100$ (quench limit).
@@
This situation contrasts the dependence on $\mathsf{H}$ of other more standard quantities in cQFTs, such as the spectrum $|\beta_{\textrm{k}}|^{2}$ of produced particles $n_{a}\sim\int{\rm d}\textrm{k}|\beta_{\textrm{k}}|^{2}/2\pi$. In general, the breakdown of parity translates into an asymmetric spectrum $|\beta_{\textrm{k}}|^{2}\neq|\beta_{-\textrm{k}}|^{2}$, which is natural in light of its effect $\hat{\psi}_{\textrm{k}}\rightarrow\hat{P}\hat{\psi}_{\textrm{k}}\hat{P}^{-1}=\gamma^{0}\hat{\psi}_{-\textrm{k}}$. However, as shown in Fig. 3, this is only manifested at intermediate values of the Hubble rate $\mathsf{H}$ in accordance with our previous results [38], but a symmetric spectrum around $\textrm{k}=0$ is recovered as one approaches the quench limit (see the darker blue curve). In the Supplemental Material [53], we indeed show how this peculiarity can be understood from the interplay of parity with the above discrete symmetries, applied to the instantaneous state after the sudden quench. We unveil an additional underlying symmetry that protects an equal particle production in modes ${\rm k}$ and $-{\rm k}$ regardless of the breaking of parity that depends on the adiabaticity parameter $\mathsf{H}$, as it is only present in the quench limit and not for slower expansions.
@@
Figure 3: Spectra of particle production for various adiabaticity regimes. The parameters are set as $ma=-1$, $g_{0}^{2}=3$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, for a chain of $N_{\rm s}=128$ sites. We highlight four values of ${\sf H}a=4$, $0.3$, $0.2$ and $0.05$ for increasingly lighter blue, for which the relation between symmetry of the spectrum and $\mathsf{H}a$ is manifest.
@@
Conclusions and outlook.– Our study highlights the potential for observing exotic analogues of gravitational particle production in effective cQFTs that go beyond the standard cosmological model. The possibility of tuning the dimensionality, designing the expansion scale factor, and tailoring the matter content of the cQFT brings a wide and largely-uncharted territory that can be addressed from a multidisciplinary viewpoint. In this work, we have focused on self-interacting Dirac QFTs in a reduced Friedman-Robertson-Walker spacetime, which have allowed us to explore how the typical phenomenon of particle-antiparticle production is affected by non-perturbative effects such as dynamical mass generation and the appearance of symmetry-broken fermion condensates. Exploiting a multidisciplinary perspective, we have analyzed this analogue cosmological setup through the lens of the entanglement contour. Making use of non-perturbative variational techniques based on fermionic Gaussian states, we have examined the spatio-temporal causal structure in the spread of entanglement carried by the produced particle-antiparticle pairs, and how the system can thermalise when looking at reduced parts. Our study reveals how the entanglement contour captures key phenomena, such as the cosmological particle horizon in a de Sitter spacetime, and how the effects of symmetry-breaking processes become very transparent when resolving the symmetries of the entanglement contour. Our work paves the way for future studies exploring further non-perturbative effects in interacting cQFTs, and encourages experimental efforts in Raman optical lattices to connect to these gravitational analogues.

### v2

@@ Abstract
We start by studying how the fermion pairs created during the expansion propagate in the asymptotically-flat out region $\mathsf{a}_{\rm f}$, where the conformal time becomes linearly related to the cosmological time $\eta=\eta_{\rm f}+(t-t_{\rm f})/\mathsf{a}_{\rm f}$ and can attain arbitrarily-large values. In Fig. 2, we depict the spatio-temporal EC profile for a partition of $\ell_{A}=128$ sites in a $N_{\rm S}=512$ lattice in real time. In the free case $g_{0}^{2}=0$, one observes a clear light-cone-like propagation of entanglement, which is consistent with the quasi-particle prediction in Eq. (11), which would be exact in the scaling limit, $\ell_{A}$, $t\to\infty$ with $t/\ell_{A}$ finite. The cones emanate from the block boundaries, separating causally connected and disconnected spacetime regions, and show a slope that is set by twice the group velocity $v_{\rm g}=\max_{\textrm{k}\in{\rm BZ}}\left|\partial_{\textrm{k}}\epsilon_{\textrm{k}}\right|$, where $\epsilon_{\textrm{k}}$ is the energy of a particle or antiparticle obtained by diagonalising $h_{\rm k}$ (2). Interestingly, this picture predicts that the entanglement equilibrates and there is an equipartition across the block $S_{i}(\eta)=S_{0}\,\,\forall i\in A$, such that the EE fulfils a volume law that agrees with the eigenstate thermalisation hypothesis [54, 55, 56] and a generalised Gibbs ensemble (GGE) that approximates $\rho_{A}$ [57].
Figure 3: (a) Evolution of entanglement entropy of a $\ell_{A}=64$ partition of a $N_{S}=512$ sites system after an expansion with a vanishing pseudo-scalar condensate $\Pi=0$. The blue line represents the numerical result while the red dashed line represents the quasi-particle prediction. The parameters are set as $ma=1$, $\mathsf{a}_{0}=0.01$, $\mathsf{a}_{\rm f}=10$, $g_{0}^{2}=3$ and ${\sf H}a=100$. (b) Evolution of the scalar and pseudo-scalar condensates for a system of $N_{S}=512$ sites, showing their persistent oscillating behaviour. The parameters are set as ma=1ma=-1, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $g_{0}^{2}=3$ and ${\sf H}a=100$. (c) Evolution of entanglement entropy of a $\ell_{A}=64$ partition of a $N_{S}=512$ sites system after an expansion with both a non-vanishing scalar and pseudo-scalar condensates, $\Sigma,\Pi\neq 0$. The blue line represents the numerical result while the red dashed line represents the quasi-particle prediction. The parameters are the same as in (b). (d) Evolution of entanglement contour of a $\ell_{A}=64$ partition of a $N_{S}=512$ sites system after an expansion with both a non-vanishing scalar and pseudo-scalar condensates, $\Sigma,\Pi\neq 0$. We show the contribution from the upper component of the spinor field to the EC, $S_{i}^{u}$. The white lines correspond to the curves $x(\eta)=\int^{\eta}v_{g}(\eta^{\prime})d\eta^{\prime}$. The parameters are the same as in (b).
@@
We now study the effect of a Gross-Neveu self-interaction (3) on the EC, considering first expansions with a vanishing pseudo-scalar condensate $\Pi(\eta)=0$ (4). In this case, the effect of interactions is to generate a scalar/chiral condensate that shifts the mass of the system dynamically $m\rightarrow m(\eta)=m+\Sigma(\eta)$. We find that, after a sudden period of exponential expansion $\mathsf{a}_{0}\to\mathsf{a}_{\rm f}$, the condensate develops some initial oscillations, after which it equilibrates to a certain non-zero value. This results in a new dispersion relation from Eq. (5) that is corrected by a time-dependent self-energy and can, consequently, lead to a dynamical group velocity $v_{\rm g}(\eta)$. In this case, the previous conserved quantities associated to the GGE become time-dependent, and may compromise the validity of the quasi-particle prediction. However, in the weakly-interacting regime, one expects to be able to derive a hydrodynamic description for the EC by renormalising the spread velocity and the particle-antiparticle mode entanglement. We confirm this expectation in the right panel of Fig. 2, where we show that the EC light cones follow $x(\eta)\approx\pm 2{v}^{\rm int}_{g}\eta$, where $v^{\rm int}_{g}$} is such a renormalised velocity. We see that the corresponding causal structure in the EC displays a compressed cone and thus a faster propagation of quantum correlations. Since the initial bare mass is negative ma=1ma=-1, a positive and non-zero scalar condensate $\Sigma(\eta)>0$ actually makes the quasi-particles lighter, leading to faster propagation. Although this is counter-intuitive at first, as the continuum QFT predicts a dynamical mass generation that would naively slow the dynamics, the problem on the lattice is richer as this dynamically-generated mass actually brings one closer to a critical line around which the continuum limit is recovered. In this case, increasing interactions reduces the gap and also the renormalised bare mass, such that the particle-antiparticle pairs propagate faster.
The initial oscillations can be understood as the results of a finite-duration quench, after which the Hamiltonian renormalised by the self-energy equilibrates becoming static, and one can find constants of motion and recover the quasi-particle picture. The agreement with the quasi-particle picture in this interacting case is shown in Fig. 3(a). We note that, due to the finite size of the system, some revivals are observed after the equilibration for $\eta\approx\frac{N_{S}}{2}$, and one has to numerically check the suitability of the quasi-particle picture by comparing the time scales of relaxation of $\Sigma(\eta)$, which is determined by the self-consistent time-evolution governed by Eq. (5), and the time scales that one is able to explore numerically for a given system size.
We now move to expansions with a non-vanishing pseudo-scalar condensate $\Pi(\eta)\neq 0$. In this case, the scalar and pseudo-scalar condensates display some synchronised and persistent oscillations that do not equilibrate for the time scales that we have explored (see Fig. 3(b)), and might connect to some long-wavelength hydrodynamic modes present at least in our fGS formalism. This leads to long-lived oscillations of the number of particles and velocities of propagation. The lack of relaxation in this case restricts the applicability of the quasi-particle picture, as shown in Fig. 3(c), where we show that any attempt to fix $s(k)$ by their later time values leads to a completely incorrect prediction. With respect to the EC, the same limitations apply, with Eq. (11) not being an accurate description of its numerical time-evolution. However, by virtue of a maximum velocity of propagation, one can still observe the presence of propagation cones in the evolution of EC, preserving the causal structure of the particle production phenomenon, as shown in Fig. 3(d). Additionally, in this case the spinor distribution of entanglement becomes non-trivial, yielding an interesting interplay with symmetry breaking as discussed in the next sections, and which is not predicted either by the quasi-particle picture.
@@
Entanglement horizons.– Let us now turn to a key phenomenon that becomes apparent via the EC dynamics during the accelerated de Sitter expansion $\mathsf{a}(\eta)$, and after it as above. For this cosmological expansion, a particle horizon determines when the distance between the emission of a particle and its subsequent detection starts to expand faster than the speed of light. Accordingly, spacetime breaks into causally disconnected regions that can have no influence on one another and, thus, share no entanglement. In conformal time, the appearance of this horizon in the EC is a bit subtle. In Fig. 4(a) we show the evolution of EC during the expansion, which shows a ballistic spread as before, but this time is upper bounded by $\eta=0$, as cosmological time squeezes in the distant future as a result of the exponential scale factor  [^1] To ensure significant propagation since the beginning of the expansion, we impose an initial state with matter content. This is achieved via a quantum quench in the mass before the start of the expansion. Other alternatives to design the initial state would lead to similar interplay with the horizon..
Figure 4: Spatio-temporal propagation of EC during a de Sitter expansion, using conformal and cosmological time respectively, and setting $ma=1$, ${\sf H}a=0.1$, $g_{0}^{2}=2$, $\mathsf{a}_{0}=1/3$, with $\eta$ and $t$ expressed in lattice units. The expansion in (a) starts at $\eta_{0}a=-30$ and extends to the distant future $\eta\to 0^{-}$, approximated by $\eta a=-0.001362$. The shaded regions denote the asymptotic spatial boundary that particles cannot cross. In (b), the corresponding expansion covers $t_{0}=0$ to $ta\to\infty$, here approximated by $t=100/a$.
@@
In Fig. 4(b), we show how resorting to cosmological time $t$ allows for a transparent account of the above particle horizon. The EC profile in cosmological time clearly shows how the propagation cones get curved as a consequence of the expansion of the universe, leading to a pair of asymptotes that separate left-right regions that cannot be causally connected. We see two separated spatial regions where entanglement attains non-vanishing values as time elapses but, due to the accelerated expansion, these are separated by a finite region with a width $\Delta x=\ell_{A}-\frac{4v_{g}}{{\sf H}\mathsf{a}_{0}}$ in which the EC is always vanishing. We emphasise that the spatial resolution brought by the EC as opposed to the EE is crucial to identify these correlation horizons. It is also interesting to note that the EC distribution within the causally-connected parts does not fulfil an equipartition, so the evolution in this case is very different from an equilibration to a GGE. Had we let the system evolve in the Minkowski out region $\mathsf{a}_{\rm f}$, the particle-antiparticle pairs would have had sufficient time to propagate in this static background, such that the causally-disconnected regions would again be connected by subsequent correlation cones as before, leading to volume-law entanglement and GGE.
@@
Symmetry breaking in the entanglement contour.– We have previously mentioned in passing that expansions in which the pseudo-scalar condensate is non-vanishing $\Pi(\eta)$ can have a bigger impact on the spread of entanglement. We now explore this further, considering that parity gets spontaneously broken, which will impact the properties of the EC under point-like transformations [28]. Specifically, the EC satisfies that $S_{i}=S_{j}$ if the system is invariant under a transformation that exchanges two sites, ${U}_{ij}\psi_{i}{U}_{ij}^{-1}=\mathord{\rm e}^{\mathord{\rm i}\xi}\psi_{j}$, where $\xi\in\mathbb{R}$ and $i\neq j$. To understand this property better, we temporarily focus on the discrete symmetries of the static model $\mathsf{a}(\eta)=1$, which connects to the tenfold classification of topological insulators [59, 60, 61]. For $\Pi=0$, the system is invariant under time-reversal T{T} and particle-hole C{C} transformations, as well as their combination S=TC{S}={T}{C}. At the single-particle level, these transformations are realised by $\mathsf{T}=T\mathcal{K}$ with $T=\gamma^{0}$, $\mathsf{C}=C\mathcal{K}$ with $C=\gamma^{0}\gamma^{1}$, and $\mathsf{S}=\gamma^{0}\gamma^{5}$, where $\mathcal{K}$ denotes complex conjugation such that $T^{\dagger}h_{-\textrm{k}}^{*}T=h_{\textrm{k}}$, $C^{\dagger}h_{-\textrm{k}}^{*}C=-h_{\textrm{k}}$ and $\mathsf{S}^{\dagger}h_{\textrm{k}}\mathsf{S}=-h_{\textrm{k}}$. These operators satisfy $\mathsf{T^{2}=C^{2}=S^{2}}=1$. In addition, parity symmetry P{P} is realised as $P=\gamma^{0}$ such that $P^{\dagger}h_{-\textrm{k}}P=h_{\textrm{k}}$. Indeed, together with translational invariance, parity symmetry describes the invariance under a mirror symmetry about any point and, in particular, about the center of the $A$ block, which leads to a site exchange ${U}_{ij}$ with $j=\ell_{A}+1-i$. According to our previous comment, if parity is conserved, one expects the EC to display the symmetry $S_{i}=S_{\ell_{A}+1-i}$, which agrees with the EC cones of Fig. 2.
@@
This situation changes when $\Pi\neq 0$, as one finds that T{T} is preserved, but P{P}, C{C} and S{S} get broken. Moreover, switching back to $\mathsf{a}(\eta)\neq 1$ will also break T{T}, endowing the condensate with its own dynamics $\Pi(\eta)$, which shall affect the EC spreading as we now discuss. In the left panel Fig. 5(a), we depict the spread of the EC in complete analogy to Fig. 2, but for an expansion with a non-zero $\Pi(\eta)$. We note that in all of the cases above, the ECs displayed actually correspond to one of the spinor components, namely $S_{i}^{\rm u}(\eta)$, which does not compromise the generality as $S_{i}^{\rm d}(\eta)=S_{i}^{\rm u}(\eta)=:S_{i}(\eta)$. This no longer holds when $\Pi(\eta)\neq 0$, and one can observe in the right panel of Fig. 5(a) that the EC for the down spinor is indeed different $S_{i}^{\rm d}(\eta)\neq S_{i}^{\rm u}(\eta)$, with a larger spread of entanglement from the left boundary with respect to the right one, opposite to what happens for $S_{i}^{\rm u}(\eta)$. These figures suggest that the EC profiles are still symmetric by simultaneously exchanging left-right boundaries and up-down spinors, yielding $S^{\textrm{u}}_{i}(\eta)=S^{\textrm{d}}_{\ell_{A}+1-i}(\eta)$ and vice versa $S^{\textrm{d}}_{i}(\eta)=S^{\textrm{u}}_{\ell_{A}+1-i}(\eta)$. In fact, one can check that this is indeed a remnant symmetry that combines both parity P{P} and particle-hole C{C} symmetries $(\gamma^{1})^{\dagger}h_{\textrm{k}}^{*}\gamma^{1}=-h_{\textrm{k}}$, which still holds even if the individual symmetries are broken. The CP{C}{P} symmetry that survives a non-zero $\Pi(\eta)$ suggests that one can define a zigzag ordering of the spatial sites and spinor components and depict the site- and spinor-resolved EC as shown in Fig. 5(b). This symmetry-aware representation allows to recover a clear light-cone-like structure characteristic of causality and connecting with the aforementioned quasi-particle picture, even in presence of a pseudo-scalar condensate and for any Hubble rate $\mathsf{H}$.
@@
Figure 5: Time evolution of EC for different phases for a partition $A$ of $N_{A}=32$ sites, from a system of $N_{S}=128$ sites. (a) Numerical time evolution of EC for each leg of the ladder, for an expansion with $\Pi\neq 0$, with ma=1ma=-1, $g^{2}=0$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $\mathsf{H}=100$ (quench limit). (b) Numerical time evolution of EC, with all the sites shown in the horizontal axis, alternating up and down spinors, for an expansion with $\Pi\neq 0$, with ma=1ma=-1, $g^{2}=0$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, $\mathsf{H}=100$ (quench limit).
@@
Let us finish our discussion by showing how the resolution offered by the EC and its interplay with the broken symmetries contrasts the dependence on the Hubble rate $\mathsf{H}$ of other more standard quantities in cQFTs, such as the spectrum $|\beta_{\textrm{k}}|^{2}$ of produced particles $n_{a}\propto\int{\rm d}\textrm{k}|\beta_{\textrm{k}}|^{2}/2\pi$. In general, the breakdown of parity translates into an asymmetric spectrum $|\beta_{\textrm{k}}|^{2}\neq|\beta_{-\textrm{k}}|^{2}$, which is natural in light of its effect $\psi_{\textrm{k}}\rightarrow{P}\psi_{\textrm{k}}{P}^{-1}=\gamma^{0}\psi_{-\textrm{k}}$. However, as shown in Fig. 6, this is only manifested at intermediate values of the Hubble rate $\mathsf{H}$ in accordance with our previous results [42], but a symmetric spectrum around $\textrm{k}=0$ is recovered as one approaches the quench limit (see the darker blue curve). In the Supplemental Material [62], we indeed show how this peculiarity can be understood from the interplay of parity with the above discrete symmetries, applied to the instantaneous state after the sudden quench. We unveil an additional underlying symmetry that protects an equal particle production in modes $\rm k$} and $\rm k$} regardless of the breaking of parity that depends on the adiabaticity parameter $\mathsf{H}$, as it is only present in the quench limit and not for slower expansions.
@@
Figure 6: Spectra of particle production for various adiabaticity regimes. The parameters are set as ma=1ma=-1, $g_{0}^{2}=3$, $\mathsf{a}_{0}=0.7$, $\mathsf{a}_{\rm f}=1.3$, for a chain of $N_{S}=128$ sites. We highlight four values of ${\sf H}a=4$, $0.3$, $0.2$ and $0.05$ for increasingly lighter blue, for which the relation between symmetry of the spectrum and $\mathsf{H}a$ is manifest.
@@
Conclusions and outlook.– Our study highlights the potential for observing exotic analogues of gravitational particle production in effective cQFTs that go beyond the standard cosmological model. The possibility of tuning the dimensionality, designing the expansion scale factor, and tailoring the matter content of the cQFT brings a wide and largely-uncharted territory that can be addressed from a multidisciplinary viewpoint. In this work, we have focused on self-interacting Dirac QFTs in a reduced Friedmann-Robertson-Walker spacetime, which have allowed us to explore how the typical phenomenon of particle-antiparticle production is affected by non-perturbative effects such as dynamical mass generation and the appearance of symmetry-broken fermion condensates. Exploiting a multidisciplinary perspective, we have analysed this analogue cosmological setup through the lens of the quasi-particle picture of entanglement entropy and the entanglement contour. Making use of non-perturbative variational techniques based on fermionic Gaussian states, we have examined the spatio-temporal causal structure in the spread of entanglement carried by the produced particle-antiparticle pairs, and how the system can thermalise when looking at reduced parts. Our study reveals how the entanglement contour captures key phenomena, such as the cosmological particle horizon in a de Sitter spacetime, and how the effects of symmetry-breaking processes become very transparent when resolving the symmetries of the entanglement contour. Our work paves the way for future studies exploring further non-perturbative effects in interacting cQFTs, and encourages experimental efforts in Raman optical lattices to connect to these gravitational analogues.